# Mapping Bias in Vision Language Models: Signposts, Pitfalls, and the Road Ahead

### v1

@@ 3 Methods
3.1 Datasets
We consider five different image datasets that include protected attributes like sex, race, religion, and age to evaluate the fairness of these models: CelebA, PATA, UTKFace, Visogender, and VLStereoSet.
3.1.1 CelebA
3.1.2 PATA
PATA is a set of about 5,000 public images crawled from the web that have been organized into 24 scenes. Each image is labeled with sex (male or female), race (White, Black, Indian, East-Asian, and Latino-Hispanic), and two age groups (youth and adults). Each image has two sets of captions associated with it: One set of positive captions with desirable descriptions of the person and one set of offensive or untoward text. These captions are developed with the protected attributes in mind and try to elicit a biased response from the model. For example, captions describing a person whose gender is the protected attribute could be a "photo of a smart engineer", trying to elicit a common bias that women cannot be smart and/or engineers. We use the entire dataset to run our evaluation as no training or test split is provided.
3.1.3 UTKFace
UTKFace is another portrait dataset that consists of over 20K portraits in the wild that cover a large variety of poses, facial expressions, illumination, etc. Each image is labeled with age, race, and sex. Similarly to Yang et al. (2024), we bin the age label into 5 categories: "Child" (020)020(0-20)( 0 - 20 ), "Young" (2040)2040(20-40)( 20 - 40 ), "Middle-Aged" (4060)4060(40-60)( 40 - 60 ), and "Senior" (60)limit-from60(60-)( 60 - ). We focus on classifying all three attributes and measure the bias based on the other two. We call each dataset using a two part name. The first part is what the model is predicting. The second part is what the bias is measured over. For example, "UTKFace Race Gender" would be predicting the race of the person while measuring the difference in performance across genders. As there is no test set provided, we sample 2,000 photos of the 20,000 as a test set to run our evaluation on.
3.1.4 VisoGender
VisoGender is a set of about 700 images that are used to benchmark gender bias in VLMs. The dataset evaluates the ability of a model to correctly resolve the pronoun of a person in a picture of them working in a professional role. The dataset measures how much the model associates different occupations with different genders. They split these images into two smaller datasets: Occupation-Object (OO) and Occupation-Participant (OP). The OO dataset has only one person working in a profession in the image, and the model must choose that person’s pronoun. The OP dataset has two people in the image: one who is working in a profession and one who is participating. The task is for the model must resolve the pronoun of the one who is working. We use the entire dataset to run our prediction as no train or test split is provided.
3.1.5 VLStereoSet
3.2 Models
We evaluate the performance and biases of various VLMs across different sizes and architectures, including both open- and closed-source models. For the open-source models, we assess the LLaVa v1.6 series, specifically: llava-v1.6-34b-hf, llava-v1.6-vicuna-7b, and llava-v1.6-mistral-7b-hf Liu et al. (2023). Additionally, we evaluate two CLIP models, CLIP L-224 and CLIP L-336, with the latter serving as the Vision Transformer (ViT) component for all the LLaVa models mentioned. For the closed-source models, we include Google’s May 2024 release: gemini-1.5-flash-001 Radford et al. (2021); Goldin et al. (2024).
@@ 3.2 Models
All experiments were conducted on Nvidia A100 80GB GPUs, with CUDA version 12.1, transformers version 4.41.0, and the API models run through Google Vertex AI with google-cloud-platform version 1.60.0. For reproducibility, we set the temperature to 0 in all cases.
Figure 2: Diagram showing the process of creating the adversarial version of the VisoGender Dataset
3.3 Prompts
The prompts used in the experiments (shown in Table 1) were adapted from those recommended by the LLaVa authors for model evaluation and inspiration The initial questions for each prompt were inspired by prompts from VQA datasets like VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), and MME Fu et al. (2024) datasets. We manually tested and refined these prompts to ensure that the models produced accurate answers on a small subset of training data.
For multiple-choice questions(MCQ), we ask the model to choose the correct caption for the image from a list of options, ending wit hrecommended prompt for MCQ type questions for LLaVa: “<MCQ Options> Answer with the option’s letter from the given choices directly.”
3.4 Bias Dataset Effectiveness Experiment
For prediction generation, we allow models to produce a distribution over the next possible tokens. From this distribution, we select the most probable token corresponding to either ("Yes", "No") for binary questions and ("A", "B", "C", etc.) for multiple-choice questions using a process similar to the lm-evaluation-harness by EleutherAI Gao et al. (2023).
To evaluate how effectively the datasets measure bias in response to visual input, we run each dataset twice: once with the image and once without the image. This approach helps us determine whether the dataset is an effective tool for measuring bias and performance in vision tasks. We hypothesize that if the dataset is indeed effective, there should be a significant performance and fairness gap between the two runs — with much better performance and fairness when the image is included.
To compare our models’ effectiveness, we include a random classifier baseline. This baseline uniformly samples an answer from the label space for each question.
3.5 Vision Reliance Experiment
To assess model robustness, we re-evaluate performance using a adversarial version of the VisoGender dataset, as shown in Figure 2. In this experiment, we modify the prompts to intentionally mislead the model by providing incorrect gender descriptions for the individuals in the image. Specifically, we append an inaccurate gender description to the end of the question. For example, if the image shows a male doctor attending to a female patient, the modified prompt might read: "Which one is the correct caption of this image of a woman helping a man?"
Despite this adversarial input, we expect the model to identify the correct caption, such as "the doctor is chatting with his patient," rather than being misled by the incorrect gender description. This experiment tests the model’s ability to rely on visual evidence rather than textual biases.
3.6 Metrics
3.6.1 Bias Dataset Effectiveness Metrics
4 Results
Figure 3: Performance of models on all the portrait datasets compared to their fairness (DPR = demographic parity ratio, higher the better.). The dashed diagonal line represents the dividing line between trading off performance for fairness or vice versa. Arrows indicate changes in performance and fairness from CLIP to LLaVa models when the differences for both are > 10%.
4.1 Portrait Datasets
Adding the LLM on top of ViT improves performance/fairness tradeoffs.
@@ 4.1 Portrait Datasets
Although integrating LLMs with CLIP models generally enhances both fairness and performance for most datasets, the tradeoffs between these two factors are inconsistent. For instance, in the CelebA blond hair dataset, we observe improvements in fairness ranging from 2 to 20 points. For UTKFace Age, we see a 20-point fairness gain when measuring performance across races. For UTKFace Race, a 4-6 point improvement across ages. In UTKFace Gender, we also see a 25-point increase in fairness when measuring across races.
However, these improvements are not universal. As mentioned, CelebA heavy makeup and UTKFace Age Gender exhibit a troubling reduction in fairness, raising the concern that fairness for one task does not reliably generalize to other tasks, even within the same image dataset. Practitioners should be cautious when assuming that fairness achieved in one domain will extend to others, even when they use the same image data.
4.2 Scene-Based Bias Evaluation
4.2.1 VLStereoSet
As shown in Figure 5 (see Appendix), the results suggest that VLStereoSet struggles to effectively evaluate whether images elicit bias in Vision Language Models (VLMs). Without the image input, all models exhibit a Vision-Language Bias Score (VLBS) around 0.5-0.6, indicating a significant level of bias since lower scores are better, and random chance would yield a score of 0.33. Including the image results in only a marginal improvement in fairness, with a bias reduction of approximately 5 points for three models and 15 points for LLaVa Vicuna 7B. This finding highlights an issue with VLStereoSet as a dataset. The modest performance improvements compared to gains observed in other datasets, such as CelebA and UTKFace, suggest that the model can often infer the correct answer without relying on the image.
@@ 4.2.1 VLStereoSet
Figure 4: Average Demographic Parity Ratio comparison between "Text Only" and "With Image" for PATA (left). Macro-F1 comparison between "Text Only" and "With Image" for PATA (right). The grey bar represents the baseline CLIP Large 336 model, which is the backbone frozen vision encoder for all LLaVa that we have tested here.
4.2.2 PATA
PATA may have some limitations when used as a bias evaluation dataset.
As seen in Figure 4, the PATA dataset also exhibits issues in its design, rendering it insufficient for evaluating bias in models. Across all models, performance either decreases or remains unchanged when images are included. This stagnation or drop in performance suggests that PATA does not require models to leverage the image input as intended.
Moreover, the fairness evaluation (Figure 4) reveals that models tend to become less fair or show no improvement when images are provided. The fact that models can achieve nearly perfect fairness using text-only inputs suggests that the inclusion of images is irrelevant and may even confuse the models. This undermines the purpose of PATA as a scene-based bias evaluation dataset, where the images should play a critical role in model assessment. Instead, PATA appears to rely solely on textual cues.
Model Dataset Acc. ↑ Average RB ↓ CF Acc. ↑
Gemini 1.5 Flash $O$ 49.56 -0.9209 0.00
OP 47.33 -0.9397 21.56
LLaVa 34B $O$ 98.68 -0.0179 80.18
OP 87.76 0.0561 57.14
LLaVa Mistral 7B $O$ 99.12 -0.0026 92.95
OP 81.41 -0.2045 58.50
LLaVa Vicuna 7B $O$ 96.48 -0.0718 62.11
OP 77.10 -0.3431 50.11
Random $O$ 50.00 0.0000 50.00
OP 50.00 0.0000 50.00
Table 2: Results for VisoGender Evaluaton. Acc is accuracy. CF Acc is the accuracy of the adversarial version of the dataset. The performance gap between Acc vs CF Acc shows how robust models are.
4.3 Pronoun Resolution with VisoGender
The original VisoGender dataset is insufficient to stress-test model bias.
In Table 2, we compare the performance of our models on both the Occupation-Object (OO) and Occupation-Participant (OP) versions of the VisoGender dataset. For the OO subset, all models, except Gemini 1.5 Flash, achieve near-perfect accuracy. On the OP dataset, the trend continues, with most models maintaining around 80% accuracy, except for Gemini 1.5 Flash, which performs worse than random. It answers 95% of the time for the female option for both datasets likely due to interference from its safety filtering mechanisms.
Fairness evaluation reveals subtle biases in the OP dataset.
For the OO subset, most models exhibit near-unbiased performance, with bias scores close to zero. However, the OP dataset reveals more nuanced biases. Specifically, the 7B LLaVa models tend to be more accurate for images where females are the primary focus, while LLaVa 34B exhibits an opposite bias, showing better accuracy for images with males.
The OP dataset provides more valuable insights into model bias than the OO subset.
When measuring the models’ ability to reveal biases, the OO dataset proves ineffective, as all LLaVa models achieve near-perfect accuracy and fairness, or, in the case of Gemini, fail to answer adequately. In contrast, the OP dataset provides more informative results: LLaVa Mistral and Vicuna models demonstrate a bias towards females, while LLaVa 34B is slightly biased toward males. These insights suggest that the OP dataset is far more useful for uncovering gender-related biases in pronoun resolution tasks than the OO version.
4.4 Adversarial Dataset for Pronoun Resolution
As shown in Table 2, model performance drops significantly when adversarial prompts are introduced, indicating a reliance on textual information even when it conflicts with visual cues. Most models experience a 20-30 point drop in accuracy from their baseline performance. Notably, LLaVa Mistral 7B shows the least degradation, with only a 6-point drop in accuracy on the OO dataset, suggesting it is more robust to adversarial prompts than the other models. On the other hand, Gemini 1.5 Flash suffers the most, with a 50-point accuracy drop on the OO dataset and a 25-point drop on the OP dataset, despite being one of the largest and most powerful models tested.
These results highlight the brittleness of models when visual and textual information conflict.
The significant accuracy drops across almost all models suggest that they tend to prioritize textual information over visual cues when the two are in conflict. This phenomenon, often referred to as "sycophancy," exposes a key weakness in current bias evaluation methods: models are highly susceptible to misleading text, which skews the bias analysis. To ensure accurate assessments of bias, datasets and prompts must be meticulously designed to avoid introducing conflicting or extraneous information.
The introduction of adversarial prompts shows that most models are less robust when the visual content is at odds with the textual description, underscoring the importance of well-designed datasets for bias evaluation. The adversarial version of VisoGender provides a much-needed stress test, revealing hidden biases that are not evident in the original dataset.
5 Discussion and Conclusion
This paper presents one of the first comprehensive evaluations of bias in VLMs, analyzing state-of-the-art models across major datasets to assess both model bias and the effectiveness of bias evaluation techniques.
‘Portrait’ datasets are effective for bias evaluation.
Portrait-based datasets like UTKFace and CelebA effectively highlight fairness gaps between LLaVa and CLIP models, showing that VLMs are generally less biased and more performant when paired with an LLM backbone. However, biases in these models vary significantly depending on the dataset, emphasizing the need for diverse datasets to capture different bias dimensions.
‘Scene’ datasets fail to accurately measure bias.
Scene-based datasets like VLStereoSet and PATA suffer from data artifacts that allow models to exploit text cues, underestimating bias. Future work could refine these datasets to challenge models to properly integrate visual and textual data.
‘Limitations of VisoGender’
Bias evaluation in VLMs is particularly challenging due to models often prioritizing textual cues over visual information, a behavior known as sycophancy. This tendency leads to biased outputs when text and images conflict. The original VisoGender dataset is no longer an effective benchmark, as models now achieve near-perfect accuracy, suggesting it does not to truly test a modern model’s ability to integrate visual information. Instead, models rely heavily on text, masking underlying biases. The adversarial version of VisoGender reveals more gender-related biases, emphasizing the potential value and need for more rigorous and carefully designed datasets that challenge models to "see" and resist being misled by textual inputs.
In conclusion, portrait datasets currently offer the best tools for bias detection, while scene-based and pronoun resolution datasets require significant improvements. Future work should focus on developing more rigorous and carefully designed datasets to ensure VLMs are both fair and reliable across a variety of tasks.
6 Limitations
We identify the following limitations in our work. First, due to the growing popularity of multimodal models, it is impractical to evaluate all available models. We have focused on a representative subset acknowledging that this might limit the generalizability of our findings. Second, some of the images for VisoGender, PATA, and VLStereoSet were unable to be retrieved due to broken hyperlinks, possibly affecting the quality of the evaluations. Third, given the limited datasets available, we recognize that the models’ fairness discussed in this paper may not fully not extend to real-world scenarios.
Acknowledgments
We would like to thank UAB Research Computing for use of their HPC infrastructure. The authors acknowledge financial support from the Google PhD Fellowship (SC), the Woods Foundation (DB, SC), the NIH (NIH-USA R01CA294033 (DB, SC), NIH-USA U54CA274516-01A1 (DB, SC), the ASTRO-ACS Clinician Scientist Development Grant ASTRO-CSDG-24-1244514 (DB). The authors also thank Google Cloud research fund for Gemini API inference costs.

### v2

@@ A.2 Prompt Table
A.3 Hardware and API Versions
@@
All experiments were conducted on Nvidia A100 80GB GPUs, with CUDA version 12.1, transformers version 4.41.0, and the API models run through Google Vertex AI with google-cloud-platform version 1.60.0.
A.4 Portrait Based Dataset Tables
Model Includes Image? Prediction Feature Macro F1 ↑ DPR ↑
LLaVa 34B Blond Hair 11.76 1.00
Heavy Makeup 28.82 1.00
Blond Hair 81.47 0.31
Heavy Makeup 75.77 0.03
LLaVa Mistral 7B Blond Hair 11.76 1.00
Heavy Makeup 28.82 1.00
Blond Hair 75.19 0.48
Heavy Makeup 57.51 0.04
LLaVa Vicuna 7B Blond Hair 46.43 $\infty$
Heavy Makeup 37.31 $\infty$
Blond Hair 82.97 0.19
Heavy Makeup 74.41 0.02
Paligemma 10B 224 Blond Hair 46.43 $\infty$
Heavy Makeup 37.31 $\infty$
Blond Hair 78.79 0.50
Heavy Makeup 60.97 0.04
Paligemma 10B 448 Blond Hair 46.43 $\infty$
Heavy Makeup 28.82 1.00
Blond Hair 73.44 0.56
Heavy Makeup 60.31 0.09
CLIP Large 224 Blond Hair 80.83 0.17
Heavy Makeup 45.31 0.62
CLIP Large 336 Blond Hair 80.25 0.20
Heavy Makeup 43.22 0.53
Table 4: Results for CelebA dataset. Predicting is the attribute that is being predicted by the model. DPR is demographic parity ratio.
Model Includes Image? Protected Category Macro F1 ↑ DPR ↑
LLaVa 34B Gender 16.51 1.00
Race 1.00
Gender 70.08 0.50
Race 0.52
LLaVa Mistral 7B Gender 8.05 $\infty$
Race $\infty$
Gender 74.02 0.49
Race 0.50
LLaVa Vicuna 7B Gender 8.76 $\infty$
Race $\infty$
Gender 73.71 0.66
Race 0.54
Paligemma 10B 224 Gender 8.76 $\infty$
Race $\infty$
Gender 28.03 0.52
Race 0.17
Paligemma 10B 448 Gender 8.76 $\infty$
Race $\infty$
Gender 36.29 0.48
Race 0.07
CLIP Large 224 Gender 51.23 0.82
Race 0.38
CLIP Large 336 Gender 51.68 0.82
Race 0.37
Table 5: Results for UTKFace Age Dataset. PC is the protected category. DPR is the demographic parity ratio.
Model Image? PC Macro F1 ↑ DPR ↑
LLaVa 34B Age 32.45 1.00
Race 1.00
Age 96.89 0.50
Race 0.71
LLaVa Mistral 7B Age 34.20 $\infty$
Race $\infty$
Age 97.05 0.51
Race 0.74
LLaVa Vicuna 7B Age 32.45 1.00
Race 1.00
Age 96.23 0.42
Race 0.69
Paligemma 10B 224 Age 34.20 $\infty$
Race $\infty$
Age 91.06 0.51
Race 0.67
Paligemma 10B 448 Age 34.20 $\infty$
Race $\infty$
Age 85.87 0.42
Race 0.58
CLIP Large 224 Age 90.50 0.51
Race 0.77
CLIP Large 336 Age 92.06 0.74
Race 0.53
Table 6: Results for UTKFace Gender Dataset. PC is the protected category. DPR is the demographic parity ratio.
Model Includes Image? Protected Category Macro F1 ↑ DPR ↑
LLaVa 34B Age 2.82 $\infty$
Gender $\infty$
Age 77.78 0.48
Gender 0.97
LLaVa Mistral 7B Age 2.82 $\infty$
Gender $\infty$
Age 71.85 0.33
Gender 0.93
LLaVa Vicuna 7B Age 11.97 $\infty$
Gender $\infty$
Age 76.76 0.45
Gender 0.90
Paligemma 10B 224 Age 11.97 $\infty$
Gender $\infty$
Age 16.73 0.48
Gender 0.41
Paligemma 10B 448 Age 11.97 $\infty$
Gender $\infty$
Age 40.58 0.27
Gender 0.91
CLIP Large 224 Age 47.53 0.41
Gender 0.80
CLIP Large 336 Age 47.68 0.34
Gender 0.83
Table 7: Results for UTKFace Race Dataset. PC is the protected category. DPR is the demographic parity ratio.
A.5 Portrait Data Analysis
Adding the LLM on top of improves performance/fairness tradeoffs.
@@ A.5 Portrait Data Analysis
Although integrating LLMs with CLIP models generally enhances both fairness and performance for most datasets, the tradeoffs between these two factors are inconsistent. For instance, in the CelebA blond hair dataset, we observe improvements in fairness ranging from 2 to 20 points. For UTKFace Age, we see a 20-point fairness gain when measuring performance across races. For UTKFace Race, a 4-6 point improvement across ages. In UTKFace Gender, we also see a 25-point increase in fairness when measuring across races. However, these improvements are not universal. As mentioned, CelebA heavy makeup and UTKFace Age Gender exhibit a troubling reduction in fairness, raising the concern that fairness for one task does not reliably generalize to other tasks, even within the same image dataset. Practitioners should be cautious when assuming that fairness achieved in one domain will extend to others, even when they use the same image data.
A.6 Scene Based Dataset Tables
Model Image? PC VLBS ↓
Gender 66.67
Profession 58.21
Race 50.00
Religion 40.00
Gender 59.69
Profession 54.89
Race 58.65
Gemini 1.5 Flash Religion 20.00
Gender 67.88
Profession 66.92
Race 61.69
Religion 20.00
Gender 59.07
Profession 45.38
Race 53.23
LLaVa 34B Religion 20.00
Gender 74.61
Profession 73.85
Race 67.16
Religion 20.00
Gender 59.16
Profession 54.20
Race 52.20
LLaVa Mistral 7B Religion 50.00
Gender 60.10
Profession 63.85
Race 51.74
Religion 20.00
Gender 40.41
Profession 41.54
Race 26.87
LLaVa Vicuna 7B Religion 20.00
Table 8: Results for VLStereoSet. PC is the protected category. VLBS is the Vision-language bias score.
Model Image? PC VLBS ↓
Gender 97.38
Profession 97.71
Race 92.20
Religion 100.00
Gender 78.53
Profession 74.05
Race 68.29
Paligemma 10B 224 Religion 66.67
Gender 97.38
Profession 97.71
Race 92.20
Religion 100.00
Gender 64.40
Profession 60.31
Race 60.98
Paligemma 10B 448 Religion 83.33
Gender 62.35
Profession 55.56
Race 64.74
CLIP Large 224 Religion 40.00
Gender 58.64
Profession 61.11
Race 58.96
CLIP Large 336 Religion 40.00
Table 9: Results for VLStereoSet. PC is the protected category. VLBS is the Vision-language bias score.
Model Includes Image? Protected Category Macro F1 ↑ DPR ↑
Age 72.85 0.77
Gender 81.53 0.97
Race 78.58 0.91
Age 61.72 0.79
Gender 85.25 0.96
Gemini 1.5 Flash Race 76.26 0.90
Age 92.54 0.98
Gender 90.98 0.97
Race 94.84 0.96
Age 95.01 0.90
Gender 86.81 0.83
LLaVa 34B Race 97.13 0.96
Age 100.00 1.00
Gender 98.41 1.00
Race 99.99 1.00
Age 92.71 0.85
Gender 92.93 0.89
LLaVa Mistral 7B Race 95.19 0.95
Age 100.00 1.00
Gender 98.40 1.00
Race 88.65 0.98
Age 93.27 0.88
Gender 95.90 0.94
LLaVa Vicuna 7B Race 98.35 0.96
Age 98.35 0.94
Gender 98.35 1.00
Race 98.35 0.99
Age 94.22 0.95
Gender 95.14 0.99
Paligemma 10B 224 Race 98.01 0.97
Age 53.90 0.77
Gender 79.50 0.97
Race 81.27 0.93
Age 92.97 0.90
Gender 90.02 0.99
Paligemma 10B 448 Race 90.36 0.96
Age 80.29 0.94
Gender 92.81 0.90
CLIP Large 224 Race 91.26 0.96
Age 80.33 0.95
Gender 92.09 0.89
CLIP Large 336 Race 91.54 0.93
Table 10: Results for PATA Dataset. PC is the protected category. DPR is the demographic parity ratio.
A.7 VLStereoSet
Figure 4: Comparison between ”Text Only” and ”With Image” of average VLBS for VLStereoSet. Grey bar represents baseline CLIP model.
A.8 PATA
@@
Figure 5: Average Demographic Parity Ratio comparison between ”Text Only” and ”With Image” for PATA (left). Macro-F1 comparison between ”Text Only” and ”With Image” for PATA (right). The grey bar represents the baseline CLIP Large 336 model, which is the backbone frozen vision encoder for all LLaVa that we have tested here.
A.9 Steering Features Tested
ID Location Index Description
layer_0/width_131k/canonical Residual 78395 The Concept of Fairness
layer_1/width_131k/canonical MLP 18813 the concept of fairness
layer_0/width_131k/canonical MLP 3358 instances of the word ”fair” and its variations, indicating discussions about fairness or equity
layer_35/width_131k/canonical Residual 67237 expressions of fairness or equity
layer_16/width_131k/canonical MLP 64818 expressions of fairness or equity
layer_18/width_131k/canonical MLP 128370 concepts and discussions surrounding gender equality and equality in general
layer_30/width_131k/canonical MLP 96507 references to gender equality and disparities
layer_35/width_131k/canonical Residual 41826 terms related to gender equity and pay disparity
layer_13/width_131k/canonical MLP 7624 terms related to gender dynamics and equity
layer_18/width_131k/canonical Residual 7958 issues related to gender equality and employment in legal and social contexts
layer_25/width_131k/canonical MLP 96024 gender-associated roles and professions
layer_26/width_131k/canonical MLP 19178 terms related to gender roles and professions
Table 11: Information about each feature steered for Paligemma-2 10B. Features were found on Neuronpedia for Gemma-2-9B. ID is the identifier for the SAE used. It contains the layer and the width of the SAE. Location is where the hidden state was taken in the layer.
A.10 Prompts Used for Debiasing
Dataset Prompt Labels
VisoGender Based on the image which caption is most contextually appropriate while considering fairness? Multiple Choice Question
Which caption aligns best with the image ensuring the choice is based on visual cues rather than assumptions?
Which caption best represents the image while avoiding gender-based assumptions?
Choose the caption that best matches the image while making a fair and unbiased decision.
Choose the caption that best matches the image while making a fair and unbiased decision remember that women can hold the same jobs as men.
A.11 Debiasing Prompting Results
Model Dataset Intervention Score
Paligemma 224 $O$ None 0.5815
Best Prompt-based Debias Technique 0.5815
Best SAE-based Debias Technique 0.6916
OP None 0.6168
Best Prompt-based Debias Technique 0.6168
Best SAE-based Debias Technique 0.6485
Adv OO None 0.5242
Best Prompt-based Debias Technique 0.5242
Best SAE-based Debias Technique 0.6784
Adv OP None 0.5011
Best Prompt-based Debias Technique 0.5011
Best SAE-based Debias Technique 0.5533
Paligemma 448 $O$ None 0.5463
Best Prompt-based Debias Technique 0.5463
Best SAE-based Debias Technique 0.7093
OP None 0.5714
Best Prompt-based Debias Technique 0.5714
Best SAE-based Debias Technique 0.6576
Adv OO None 0.5066
Best Prompt-based Debias Technique 0.5066
Best SAE-based Debias Technique 0.6211
Adv OP None 0.4898
Best Prompt-based Debias Technique 0.4898
Best SAE-based Debias Technique 0.5351
Table 12: Results for debiasing evaluation on Visogender ($O$, OP) and Adversarial Visogender (Adv $O$, Adv OP). Acc is accuracy. Best debiasing and prompt method chosen for each dataset and model. Avg. RB is average resolution bias.